# MH-MoE: Multi-Head Mixture-of-Experts

### v1

@@ Abstract
Multi-Head Mixture-of-Experts (MH-MoE) [16] demonstrates superior performance by using the multi-head mechanism to collectively attend to information from various representation spaces within different experts. In this paper, we present a novel implementation of MH-MoE that maintains both FLOPs and parameter parity with sparse Mixture of Experts models. Experimental results on language models show that the new implementation yields quality improvements over both vanilla MoE and fine-grained MoE models. Additionally, our experiments demonstrate that MH-MoE is compatible with 1-bit Large Language Models (LLMs) such as BitNet [9].
@@ 3 Experiments Setup
4 Experiments
5 Language Modeling Evaluation
@@
SMoE 100,000 10.41 9.44 14.30
Fine-grained SMoE 10.66 9.15 13.78

### v3

@@ Abstract
Multi-Head Mixture-of-Experts (MH-MoE) [16] demonstrates superior performance by using the multi-head mechanism to collectively attend to information from various representation spaces within different experts. In this paper, we present a novel implementation of MH-MoE that maintains both FLOPs and parameter parity with sparse Mixture of Experts models. Experimental results on language modeling tasks indicate that the new implementation yields quality improvements over both vanilla MoE and fine-grained MoE models. Additionally, our experiments show that MH-MoE is compatible with 1-bit Large Language Models (LLMs) such as BitNet [9].
@@ 3 Experiments
3.1 Language Modeling Evaluation
@@
SMoE 100,000 10.66 9.44 14.30
Fine-grained SMoE 10.41 9.15 13.78